OpenStudy (anonymous):

Calculus Question, homies <3 If a ball is thrown vertically upward from the roof of a 32 foot tall building with a velocity of 16 ft/sec, its height after t seconds is s(t) = 32 + 16 t - 16 t^2. What is the maximum height the ball reaches? What is the velocity of the ball when it hits the ground (height 0)?

geerky42 (geerky42):

1) You may want to take derivative then set it equal to 0 then solve for t. We can call this t "\(t_{top}\)" 2) You would need to find \(\left.\dfrac{ds(t)}{dt}\right|_{~t=2t_{top}}=v(2t_{top})\) Since it take \(t_{top}\) for ball to reaches maximum height and another \(t_{top}\) for ball to hit ground, hence \(2t_{top}\) Does that make sense?

OpenStudy (anonymous):

Sorry, I was in between classes. So the derivative of what you were saying is v(2ttop)? Where does the 32 feet come in to play?

OpenStudy (anonymous):

If I take the derivative of that equation, it would be -32t+16?

OpenStudy (anonymous):

I was able to figure this one out. First, I got the derivative of s(t)=16t-32t^2+32. so s'(t)=16-32t. I solved for t and ended up getting .5. I put .5 in to the equation 16t-32t^2+32 and ended up getting 36 for the answer For the second answer, I set 16t-32t^2+32 equal to 0 and solved. I ended up getting -1 and 2. I plug 2 into the derivative and got -48!